Red laser light from a He-Ne laser (λ = 632.8 nm) creates a second-order fringe at 53.2° after passing through the grating. What is the wavelength λ of light that creates a first-order fringe at 18.8°?
Answer #1

given,

wavelength of He-Ne laser = 632.8 nm

creates second order fringe at 53.2 degree

sin(theta) = m * wavelength / d

sin(53.2) = 2 * 632.8 * 10^-9 / d

d = 1.58056×10^-6 m

in other case

sin(18.8) = 1 * wavelength / (1.58056×10^-6)

wavelength = 5.0936 × 10^-7 m